### v1

@@ Appendix B Proofs of Theorems
Denote $r_{\lambda}=(r_{1\lambda},\cdots,r_{n\lambda})^{\prime}$, $r_{\lambda\delta}=r_{\lambda}+r_{\delta}$. Thus, $\hat{\vartheta}=\vartheta+R\hat{\xi}=\vartheta+R(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}{P_{K_{3}}}(r_{\lambda\delta}+Bv)$ and $\big{\|}r_{\lambda\delta}\big{\|}^{2}=o(1)$ by Assumption 16. Under $H_{0}:\vartheta=0$,
$\begin{split}n\hat{\alpha}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\hat{\alpha}=(v^{\prime}\widehat{\mathcal{M}}_{4}v+A_{11}+A_{12}+2A_{21}+2A_{22})\end{split}$ (B.13)
where $A_{11}=r_{\lambda\delta}^{\prime}(\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4})r_{\lambda\delta},A_{12}=r_{\lambda\delta}^{\prime}\mathscr{M}_{4}r_{\lambda\delta}$, $A_{21}=r_{\lambda\delta}^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v$, $A_{22}=r_{\lambda\delta}^{\prime}\mathcal{M}_{4}v$, ^$\widehat{\mathcal{M}}_{4}=\frac{1}{n}B^{\prime}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}B$ with $\widehat{\mathcal{V}}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\widehat{\mathscr{D}}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$, $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$, $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\allowbreak\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}$, and ^$\widehat{\mathscr{M}}_{4}=\frac{1}{n}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}$. Observe that $\mathcal{U}_{3}=T_{3}^{-1}\left(\frac{1}{n}L^{\prime}K_{3}\right)\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\Xi_{3}\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\left(\frac{1}{n}K_{3}^{\prime}L\right)T_{3}^{-1}$, we have $\big{\|}\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big{\|}\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ according to the proof of Lemma 5.1 and Assumptions 15 and 17. Thus, $\big{\|}\widehat{\mathcal{U}}_{3}^{-1}-\mathcal{U}_{3}^{-1}\big{\|}\leq\big{\|}\widehat{\mathcal{U}}_{3}^{-1}\big{\|}\big{\|}\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big{\|}\big{\|}\mathcal{U}_{3}^{-1}\big{\|}\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ by Assumptions 6 and 17, where $r_{n}=\frac{\ell d_{\gamma}^{2}}{n}+\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}$. Similarly, we have ^$\big{\|}\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4}\big{\|}=o_{p}(1)$, and ^$\big{\|}\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4}\big{\|}=o_{p}(1)$ by the structure of $\mathscr{M}_{4}$, $\mathcal{M}_{4}$ and Assumptions 7 and 15.
Thus, we have
@@
since $\mathrm{E}(v_{i}^{2})=1$.
Similar to (B.14) and the proof of Theorem 2.2, we have $\frac{1}{\sqrt{d_{\alpha}}}\left|A_{11}\right|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$, $\frac{1}{\sqrt{d_{\alpha}}}|A_{12}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ since $\mathscr{M}_{4}=O_{p}(1)$ by Assumptions 7 and 15, $\frac{1}{\sqrt{d_{\alpha}}}|A_{21}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ and $\frac{1}{\sqrt{d_{\alpha}}}|A_{22}|=o_{p}(1)$. Therefore, we obtain the sufficient rate from (5.7) as $n\rightarrow\infty$ for the required quantities to be negligible.
@@
(i) Set $J=J_{3}$, $\mathscr{A}=\mathcal{M}_{4}$ and $\mathscr{S}=\frac{1}{n}K_{3}\mathcal{V}_{4}K_{3}^{\prime}$ in Theorem A.1. We now proceed as in the proof Theorem 4.3 of Gupta and Qu (2024) and note that the entries of $\mathscr{S}$ are,
$s_{ij,4}=\frac{1}{n}k_{i}^{\prime}\mathcal{V}_{4}k_{j},$ (B.15)
where $k_{i}^{\prime}$ is the $i$-th row of $K_{3}$, and so
$\left|s_{ij,4}\right|=O_{p}\left(n^{-1}\left\|k_{i}\right\|\left\|k_{j}\right\|\right)=O_{p}\left(d_{\gamma}n^{-1}\right),$ (B.16)
uniformly in $i,j$, by Assumptions 15 and 17, and arguments used in the proof of Theorem 5.1 above.
Similarly, we also observe that
$\sum_{j=1}^{n}s_{ij,4}^{2}=O_{p}\left(d_{\gamma}n^{-1}\right),$ (B.17)
uniformly in $i$. Thus Assumption A(b) is satisfied. Assumption A is satisfied to due to Assumption 17. The conclusion now follows by Theorem A.1.
(ii) This proof is similar to that of Theorem 4.4 in Gupta and Qu (2024).
@@
(iii) We prove all local alternatives here. We define $H_{\ell i}\equiv H_{\ell i,n}$, with the corresponding statistic being $\mathbb{W}_{i}^{*}$, for $i=1,2,3,4$. For $\mathbb{W}_{1}^{*}$, we consider $\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$ with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix. Under $H_{\ell 1}$, we have
$\mathbb{W}_{1}^{*}=\mathbb{W}_{1}+\frac{1}{\sqrt{2d_{\alpha}}}d_{\alpha}^{\frac{1}{2}}\mathcal{A}_{1}(\frac{1}{n}\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{-1}+o_{p}(1),$
where $\mathcal{A}_{1}=\frac{1}{n}\nu_{1n}^{\prime}\mathscr{D}_{1}^{-1}\nu_{1n}$. We choose $\Gamma_{1n}=\mathscr{D}_{1}^{-1}$ and then finish the proof. For $\mathbb{W}_{4}^{*}$, we consider $\vartheta=\vartheta^{*}\equiv\nu_{4n}d_{\vartheta}^{\frac{1}{4}}/(n\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n})^{\frac{1}{2}}$. Under $H_{\ell 4}$, we have
$\mathbb{W}_{4}^{*}=\mathbb{W}_{4}+\frac{1}{\sqrt{2d_{\vartheta}}}d_{\vartheta}^{\frac{1}{2}}\mathcal{A}_{4}(\frac{1}{n}\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n})^{-1}+o_{p}(1),$
where $\mathcal{A}_{4}=\frac{1}{n}\nu_{4n}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\nu_{4n}$, and $\big{\|}\frac{1}{n}\nu_{4n}^{\prime}(\widehat{\mathscr{D}}_{4}^{-1}-\mathscr{D}_{4}^{-1})\nu_{4n}\big{\|}=o_{p}(1)$ by (5.7). We choose $\Gamma_{4n}=\widehat{\mathscr{D}}_{4}^{-1}$ and then finish the proof. The proofs of $\mathbb{W}_{2}^{*}$ and $\mathbb{W}_{3}^{*}$ are similar. ∎
@@ Appendix C Empirical results
Table 1: Data descriptions in Section 6.1, logarithm form.
Variables Descriptions Mean Min Max Sd
$y$ Value added 11.539 3.611 19.216 1.847
$p_{1}$ Capital assets 10.280 1.609 19.160 1.980
$p_{2}$ Average number of employ 5.647 2.021 11.865 1.130
$z$ Management expense 7.707 2.303 16.139 1.853
Table 2: Results of baseline model according to (6.1) in Section 6.1.
                 $\delta_{0}$                  3.313***
                 $\delta_{1}$                  0.554***
                 $\delta_{2}$                  0.402***
                 $\beta_{1}$                  0.224
                 $\beta_{2}$                  0.018
                 $H_{0}:\delta_{1}+\delta_{2}-1=0$
                 t-test                  -0.238
                 p-value                  0.812
                 $n$                  7355
                 The standard errors of the estimators
                 are reported in parentheses.
                  $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 3: Estimation and testing results in Section 6.1 using polynomial series with $h=2$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
@@
$\beta_{0}$ 1.073** 0.167* 0.153** 0.167*** 0.153*** -0.485*** -0.093*** 0.077***
(0.514) (0.100) (0.076) (0.016) (0.020) (0.013) (0.004) (0.003)
$\alpha_{11}$ 1.254** -0.036*** -0.032*** -0.036 -0.032 -0.023 0.005 0.004
(0.551) (0.007) (0.004) (0.027) (0.035) (0.975) (0.033) (0.026)
$\alpha_{12}$ 0.375 -0.015 0.017 -0.015 0.017 -0.032 0.015*** 0.027***
(3.048) (0.035) (0.019) (0.010) (0.036) (0.050) (0.002) (0.002)
$\alpha_{21}$0.194 -0.004 -0.004 -0.004 -0.004 -0.002 0.000 0.003
(0.786) (0.020) (0.033) (0.005) (0.012) (0.111) (0.001) (0.002)
$\alpha_{22}$0.023 0.092*** 0.088*** 0.092*** 0.088*** 0.032*** -0.017*** 0.012***
(0.077) (0.013) (0.013) (0.004) (0.029) (0.008) (0.000) (0.000)
$\alpha_{31}$0.046 0.031 0.019 0.031 0.019* 0.011 0.001 0.003
(0.891) (0.017) (0.010) (0.022) (0.011) (0.136) (0.018) (0.014)
$\alpha_{32}$0.062 0.004 0.005 0.004 0.005 0.059 -0.033 0.026
(0.092) (0.023) (0.023) (0.001) (0.004) (0.060) (0.031) (0.051)
@@
$\mathscr{T}_{n}$ 0.451 0.191 0.194 0.492 0.134 0.893 0.424 0.274
$\mathscr{T}_{n}^{*}$ 0.767 0.213 0.167 0.413 0.167 0.867 0.433 0.213
$H_{0}:\mu=0$ 68.365 92.106
$\mathscr{T}_{n}$ $<$0.001 $<$0.001
$\mathscr{T}_{n}^{*}$ 0.007 0.007
$n$ 7355
The standard errors of estimators are reported in parentheses. $p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 4: Estimation and testing results in Section 6.1 using trigonometric series with $h=2$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
@@
$\beta_{0}$ 1.111*** -0.381*** -0.382** -0.381*** -0.382*** -0.187*** 0.747*** 1.076***
(0.020) (0.136) (0.162) (0.042) (0.021) (0.012) (0.245) (0.380)
$\alpha_{11}$ 0.949*** 0.167*** 0.160*** 0.167*** 0.160*** 0.125*** -0.253*** -0.024***
(0.150) (0.023) (0.025) (0.014) (0.012) (0.016) (0.025) (0.004)
$\alpha_{12}$ 0.861*** -0.049 -0.028** -0.049 -0.028 -0.124*** 0.146*** 0.034***
(0.151) (0.033) (0.012) (0.035) (0.032) (0.008) (0.040) (0.007)
$\alpha_{21}$ 0.001 0.0096 0.094 0.096*** 0.094*** 0.014*** -0.091*** 0.021***
(0.007) (0.112) (0.124) (0.003) (0.003) (0.002) (0.019) (0.005)
$\alpha_{22}$ 0.002 0.026 0.024 0.026*** 0.024*** -0.018*** 0.020*** 0.009***
(0.007) (0.069) (0.043) (0.004) (0.003) (0.002) (0.005) (0.003)
$\alpha_{31}$ 0.017 0.085 0.068 0.085 0.068 0.001 -0.020 0.019
(0.018) (0.246) (0.191) (0.105) (0.064) (0.001) (0.017) (0.012)
$\alpha_{32}$ 0.009 0.033 0.025 0.033 0.025 -0.003 0.047 0.021
(0.022) (0.090) (0.084) (0.002) (0.002) (0.003) (0.007) (0.006)
@@
$\mathscr{T}_{n}$ 0.678 0.555 0.645 0.581 0.624 0.455 0.426 0.338
$\mathscr{T}_{n}^{*}$ 0.660 0.533 0.680 0.553 0.600 0.467 0.267 0.333
$H_{0}:\mu=0$ 321.305 406.511
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.013 0.027
$n$ 7355
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 5: Estimation and testing results in Section 6.1 using polynomial series with $h=4$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$
@@
$\beta_{0}$ 1.866 0.040*** 0.005*** 0.040*** 0.005 -0.036 0.075 -0.278***
(4.659) (0.011) (0.001) (0.011) (0.023) (0.159) (0.221) (0.062)
$\alpha_{11}$ 1.471 0.002*** 0.000*** 0.002 0.000 0.000 0.003 -0.006
(4.108) (0.000) (0.000) (1.759) (2.686) (10.471) (26.990) (3.119)
$\alpha_{12}$1.067 0.005*** 0.002*** 0.005 0.002 0.010 0.009 -0.005
(5.407) (0.001) (0.000) (0.329) (0.588) (2.398) (5.235) (0.903)
$\alpha_{13}$ 1.186 0.013*** 0.007*** 0.013 0.007 0.035 0.014 0.010
(8.484) (0.001) (0.002) (0.028) (0.053) (0.232) (0.466) (0.100)
$\alpha_{14}$0.020 0.121*** -0.004 0.121*** -0.004*** -0.174*** 0.180*** -0.046***
(0.288) (0.021) (0.042) (0.001) (0.002) (0.008) (0.016) (0.004)
$\alpha_{21}$0.032 0.015*** 0.003*** 0.015 0.003 0.000 0.006 -0.004
(0.676) (0.004) (0.001) (0.180) (0.462) (0.861) (3.899) (0.134)
$\alpha_{22}$0.029 0.041*** 0.020*** 0.041 0.020 0.004 0.013 0.022
(2.100) (0.008) (0.006) (0.031) (0.100) (0.190) (0.642) (0.071)
$\alpha_{23}$ 0.003 -0.010 0.092*** -0.010*** 0.092*** -0.011 -0.018 0.019*
(0.213) (0.019) (0.026) (0.002) (0.009) (0.017) (0.046) (0.010)
$\alpha_{24}$ $0$0.005*** -0.003** -0.005*** -0.003*** 0.005*** -0.001 -0.001***
(0.007) (0.000) (0.001) (0.000) (0.000) (0.001) (0.001) (0.000)
$\alpha_{31}$ 0.005 -0.005 -0.002*** -0.005 -0.002 -0.001 -0.002 0.006
(0.012) (0.061) (0.001) (0.344) (0.556) (1.561) (3.200) (0.827)
$\alpha_{32}$0.009 -0.008 -0.009 -0.008 -0.009 -0.009 -0.002 0.004
(0.042) (0.042) (0.053) (0.068) (0.124) (0.428) (0.543) (0.297)
$\alpha_{33}$ 0.007 0.059 -0.041 0.059 -0.041 -0.032 0.029 0.031
(0.032) (0.096) (0.113) (0.046) (0.033) (0.041) (0.040) (0.031)
$\alpha_{34}$0.002 0.001 0.016*** 0.001 0.016 0.008 -0.004 -0.003
(0.026) (0.073) (0.157) (0.006) (0.062) (0.011) (0.083) (0.044)
@@
$\mathscr{T}_{n}$ 0.963 0.426 0.481 0.301 0.405 0.985 0.952 0.825
$\mathscr{T}_{n}^{*}$ 0.947 0.440 0.500 0.287 0.467 0.933 0.887 0.793
$H_{0}:\mu=0$ 101.231 122.398
$\mathscr{T}_{n}$ $<0.01$ $<0.01$
$\mathscr{T}_{n}^{*}$ 0.007 0.007
$n$ 7355
The standard errors of estimators are reported in parentheses.
$p<0.1$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
Table 6: Estimation and testing results in Section 6.1 using trigonometric series with $h=4$.
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{W}$ $W_{n}^{p}$ $W_{n}^{d}$

### v2

@@ Appendix B Proofs of Theorems
Denote $r_{\lambda}=(r_{1\lambda},\cdots,r_{n\lambda})^{\prime}$, $r_{\lambda\delta}=r_{\lambda}+r_{\delta}$. Thus, $\hat{\vartheta}=\vartheta+R\hat{\xi}=\vartheta+R(G^{\prime}{P_{K_{3}}}G)^{-1}G^{\prime}{P_{K_{3}}}(r_{\lambda\delta}+Bv)$ and $\big\|r_{\lambda\delta}\big\|^{2}=o_{p}(1)$ by Assumption 16. Under $H_{0}:\vartheta=0$,
$\begin{split}n\hat{\vartheta}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\hat{\vartheta}=(v^{\prime}\widehat{\mathcal{M}}_{4}v+A_{11}+A_{12}+2A_{21}+2A_{22})\end{split}$ (B.12)
where $A_{11}=r_{\lambda\delta}^{\prime}(\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4})r_{\lambda\delta},A_{12}=r_{\lambda\delta}^{\prime}\mathscr{M}_{4}r_{\lambda\delta}$, $A_{21}=r_{\lambda\delta}^{\prime}(\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4})v$, $A_{22}=r_{\lambda\delta}^{\prime}\mathcal{M}_{4}v$, $\widehat{\mathcal{M}}_{4}=\frac{1}{n}B^{\prime}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}B$, $\widehat{\mathcal{V}}_{4}=(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}R^{\prime}\widehat{\mathscr{D}}_{4}^{-1}RT_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}$, $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$, $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\allowbreak\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}GT_{3}^{-1}$, and $\widehat{\mathscr{M}}_{4}=\frac{1}{n}K_{3}\widehat{\mathcal{V}}_{4}K_{3}^{\prime}$. Observe that $\mathcal{U}_{3}=T_{3}^{-1}\left(\frac{1}{n}L^{\prime}K_{3}\right)\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\Xi_{3}\left(\frac{1}{n}K_{3}^{\prime}K_{3}\right)^{-1}\left(\frac{1}{n}K_{3}^{\prime}L\right)T_{3}^{-1}$, we have $\big\|\widehat{\mathcal{U}}_{3}-\mathcal{U}_{3}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ according to the proof of Lemma 5.1 and Assumptions 15 and 17. Thus, $\big\|\widehat{\mathcal{U}}_{3}^{-1}-\mathcal{U}_{3}^{-1}\big\|\leq\big\|\widehat{\mathcal{U}}_{3}^{-1}\big\|\big\|\mathcal{U}_{3}-\widehat{\mathcal{U}}_{3}\big\|\big\|\mathcal{U}_{3}^{-1}\big\|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ by Assumptions 6 and 17, where $r_{n}=\frac{\ell d_{\gamma}^{2}}{n}+\frac{\ell^{1-1/q}d_{\gamma}}{n^{1/2-1/q}}$. Similarly, we have $\big\|\widehat{\mathscr{M}}_{4}-\mathscr{M}_{4}\big\|=o_{p}(1)$, and $\big\|\widehat{\mathcal{M}}_{4}-\mathcal{M}_{4}\big\|=o_{p}(1)$ by the structure of $\mathscr{M}_{4}$, $\mathcal{M}_{4}$ and Assumptions 7 and 15.
Since $\mathrm{E}(v_{i}^{2})=1$, we have
@@
Similar to (B.13) and the proof of Theorem 2.2, we have $\frac{1}{\sqrt{d_{\vartheta}}}\left|A_{11}\right|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{12}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ since $\mathscr{M}_{4}=O_{p}(1)$ by Assumptions 7 and 15, $\frac{1}{\sqrt{d_{\vartheta}}}|A_{21}|\leq Cr_{n}\cdot O_{p}\left(1\right)=o_{p}(1)$ and $\frac{1}{\sqrt{d_{\vartheta}}}|A_{22}|=o_{p}(1)$. Therefore, we obtain the sufficient rate from (5.7) as $n\rightarrow\infty$ for the required quantities to be negligible. ∎
@@
(i) Set $J=J_{3}$, $\mathscr{A}=\mathcal{M}_{4}$ and $\mathscr{S}=\frac{1}{n}K_{3}\mathcal{V}_{4}K_{3}^{\prime}$ in Theorem A.1. We now note that the entries of $\mathscr{S}$ are $s_{ij,4}=\frac{1}{n}k_{i}^{\prime}\mathcal{V}_{4}k_{j},$ where $k_{i}^{\prime}$ is the $i$-th row of $K_{3}$, and so $\left|s_{ij,4}\right|=O_{p}\left(n^{-1}\left\|k_{i}\right\|\left\|k_{j}\right\|\right)=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i,j$, by Assumptions 15 and 17, and arguments used in the proof of Theorem 5.1 above. Similarly, we also observe that $\sum_{j=1}^{n}s_{ij,4}^{2}=O_{p}\left(d_{\gamma}n^{-1}\right),$ uniformly in $i$. Thus Assumption A(b) is satisfied. Assumption A is satisfied to due to Assumption 17. The conclusion now follows by Theorem A.1.
(ii) By Lemma 5.1, we have
$\left\|\hat{\vartheta}-\vartheta\right\|=O_{p}\left(\sqrt{d_{\vartheta}/n}\right).$ (B.14)
From Assumptions 15 and 17, we have $\left\|\widehat{\mathscr{D}}_{4}-\mathscr{D}_{4}\right\|=o_{p}(1)$ and
$\left\|\widehat{\mathscr{D}}_{4}^{-1}-\mathscr{D}_{4}^{-1}\right\|\leq\left\|\widehat{\mathscr{D}}_{4}^{-1}\right\|\left\|\mathscr{D}_{4}-\widehat{\mathscr{D}}_{4}\right\|\left\|\mathscr{D}_{4}^{-1}\right\|=o_{p}(1).$ (B.15)
Write $\hat{\vartheta}=\vartheta+\Delta_{\vartheta}$ from (B.14) where $\Delta_{\vartheta}=O_{p}\left(\sqrt{d_{\vartheta}/n}\right)$. Then $n\hat{\vartheta}^{\prime}\hat{\mathscr{D}}_{4}^{-1}\hat{\vartheta}=n\vartheta^{\prime}\mathscr{D}_{4}^{-1}\vartheta+2n\vartheta^{\prime}\mathscr{D}_{4}^{-1}\Delta_{\vartheta}+n\Delta_{\vartheta}^{\prime}\mathscr{D}_{4}^{-1}\Delta_{\vartheta}+nR_{n},$ where $\left|R_{n}\right|=\left|\hat{\vartheta}^{\prime}\hat{\mathscr{D}}_{4}^{-1}\hat{\vartheta}-\hat{\vartheta}^{\prime}\mathscr{D}_{4}^{-1}\hat{\vartheta}\right|=o_{p}(1)\cdot\left\|\hat{\vartheta}\right\|^{2}=o_{p}\left(\|\vartheta\|^{2}+d_{\vartheta}/n\right)=o_{p}\left(\|\vartheta\|^{2}\right)$ by (B.15).
Using (5.8), (B.14) and bounded eigenvalues in (B.15),
$\begin{split}&\frac{1}{\sqrt{2d_{\vartheta}}}2n\vartheta^{\prime}\mathscr{D}_{4}^{-1}\Delta_{\vartheta}=O\left(\frac{n}{\sqrt{d_{\vartheta}}}\right)O\left(\|\vartheta\|\right)O_{p}(1)O_{p}\left(\sqrt{d_{\vartheta}/n}\right)=o_{p}\left(\frac{n\|\vartheta\|}{\sqrt{d_{\vartheta}}}\right)\\ &\frac{1}{\sqrt{2d_{\vartheta}}}n\Delta_{\vartheta}^{\prime}\mathscr{D}_{4}^{-1}\Delta_{\vartheta}=O\left(\frac{n}{\sqrt{d_{\vartheta}}}\right)O_{p}(1)O_{p}\left(\frac{d_{\vartheta}}{n}\right)=o_{p}\left(\frac{n}{\sqrt{d_{\vartheta}}}\right)\\ &\frac{1}{\sqrt{2d_{\vartheta}}}nR_{n}=o_{p}\left(\frac{n\|\vartheta\|^{2}}{\sqrt{d_{\vartheta}}}\right).\end{split}$ (B.16)
Note that for the sparse case, $\|\vartheta\|=O(1)$, and for the dense case, $\|\vartheta\|_{\infty}=O(1)$ and $\|\vartheta\|=\sqrt{d_{\vartheta}}$. Thus, in general, $\mathbb{W}_{4}=\frac{n\vartheta^{\prime}\mathscr{D}_{4}^{-1}\vartheta-d_{\vartheta}}{\sqrt{2d_{\vartheta}}}+o_{p}\Big(\frac{n\|\vartheta\|^{2}}{\sqrt{d_{\vartheta}}}\Big)$. For any nonstochastic sequence $\{C_{n}\},C_{n}=o\left(\frac{n\|\vartheta\|^{2}}{\sqrt{d_{\vartheta}}}\right),\lim_{n\to\infty}P\left(\left|\mathbb{W}_{4}\right|>C_{n}\right)=1$, so that consistency follows.
@@
(iii) We cover all local alternatives here. We define $H_{\ell i}\equiv H_{\ell i,n}$, with corresponding statistic $\mathbb{W}_{i}^{*}$, for $i=1,2,3,4$. For $\mathbb{W}_{1}^{*}$, we consider $\alpha=\alpha^{*}\equiv\nu_{1n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n})^{\frac{1}{2}}$ with $\nu_{1n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{1n}$ a $d_{\alpha}\times d_{\alpha}$ matrix. Under $H_{\ell 1}$, we have $\mathbb{W}_{1}^{*}=\mathbb{W}_{1}+\frac{1}{\sqrt{2d_{\alpha}}}d_{\alpha}^{\frac{1}{2}}\mathcal{A}_{1}\left(\frac{1}{n}\nu_{1n}^{\prime}\Gamma_{1n}\nu_{1n}\right)^{-1}+o_{p}(1),$ where $\mathcal{A}_{1}=\frac{1}{n}\nu_{1n}^{\prime}\mathscr{D}_{1}^{-1}\nu_{1n}$. We choose $\Gamma_{1n}=\mathscr{D}_{1}^{-1}$ and then finish the proof. For $\mathbb{W}_{4}^{*}$, we consider $\vartheta=\vartheta^{*}\equiv\nu_{4n}d_{\vartheta}^{\frac{1}{4}}/\left(n\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n}\right)^{\frac{1}{2}}$. Under $H_{\ell 4}$, we have $\mathbb{W}_{4}^{*}=\mathbb{W}_{4}+\frac{1}{\sqrt{2d_{\vartheta}}}d_{\vartheta}^{\frac{1}{2}}\mathcal{A}_{4}\left(\frac{1}{n}\nu_{4n}^{\prime}\Gamma_{4n}\nu_{4n}\right)^{-1}+o_{p}(1),$ where $\mathcal{A}_{4}=\frac{1}{n}\nu_{4n}^{\prime}\widehat{\mathscr{D}}_{4}^{-1}\nu_{4n}$, and $\left\|\frac{1}{n}\nu_{4n}^{\prime}(\widehat{\mathscr{D}}_{4}^{-1}-\mathscr{D}_{4}^{-1})\nu_{4n}\right\|=o_{p}(1)$ by (5.7). We choose $\Gamma_{4n}=\widehat{\mathscr{D}}_{4}^{-1}$ and then finish the proof. The proofs of $\mathbb{W}_{2}^{*}$ and $\mathbb{W}_{3}^{*}$ are similar. ∎
@@ Appendix C Empirical Results
Table 1: Estimation and testing results in Section 6.1.
Panel A: polynomial series, $h=2$
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
@@
$H_{0}:\mu=0$ 68.365*** 92.106***
Panel B: trigonometric series, $h=2$
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
@@
$H_{0}:\mu=0$ 321.305*** 406.511***
Panel C: polynomial series, $h=4$
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$
@@
$H_{0}:\mu=0$ 101.231*** 122.398***
Panel D: trigonometric series, $h=4$
Matrices $W_{n}^{p}$ $W_{n}^{d}$ $W_{n}^{p}$ $W_{n}^{d}$ $\widehat{G}$ $W_{n}^{p}$ $W_{n}^{d}$